# DySLIM: Dynamics Stable Learning by Invariant Measure for Chaotic Systems

### v1

@@ Appendix F Further Results
F.1 Lorenz 63 System
Figure 4 depicts the attractor for the Lorenz $63$ for a ensemble of trajectories (with randomly chosen initial conditions close to the attractor) at time $t=400$. We can observe that the regularization provides some extra symmetry to the attractor, as observed by a better defined right-wing.
In addition, we provide several other metrics to showcase the advantage of our methodology. Figure 5 shows the behavior of the MMD metric for much longer horizons to the ones used for training. We can observe that DySLIM outperforms all the baselines. We also, studied the Wasserstein-1 distance of different features involved in the ground truth dynamics shown in Equation 32. In this case Figure 6 shows that the distribution of each of the components is better captured in the models trained with DySLIM.
F.2 Kuramoto-Sivashinsky System
We provide extra numerical evidence on the stability of the trajectories generated by training the ML-model using the proposed regularization. Besides the results shown in the main text, Figure 7 additionally shows this property by comparing the distribution of first order (${\bm{u}}_{x}$) and second order (${\bm{u}}_{xx}$) spatial derivatives for ground truth and predicted trajectories, which are relevant quantities that appear in the PDE that defines this system in Equation 33. We use finite difference methods to compute ${\bm{u}}_{x}$ and ${\bm{u}}_{xx}$ at each point in time for each trajectory in the test set. At each point in time, we thus have a distribution for these derivatives across test set trajectories and spatial grid. We use the Wasserstein-1 distance to compare ground truth and predicted distributions and find that models trained with regularized objectives better match the distribution of ground truth spatial derivatives. In summary, Figure 7 shows that by regularizing the loss, we obtain a closer distribution on the derivatives than when using the unregularized loss. In fact, for some cases of the curriculum training, the Wasserstein-1 matrix explodes as the trajectories are highly unstable.
Figure 7: Wasserstein distance metric on distribution of first order (top) and second order (bottom) spatial derivatives across time. Values greater than 3 are not shown on the plot. Each line corresponds to one of five random training seeds with bolded lines indicating median values (excluding trajectories that produce NaN values).
@@ F.3 Kolmogorov Flow
In this section we provide further samples as the one shown in the main body of the paper, including ablation results.
Table 2 compiles extra results for the model trained using curriculum training with and without regularization following the description in Appendix E.3.1. Table 2 shows that using DySLIM effectively decreased, or remained competitive, the error in all the considered metrics. We point out that for Curr training the error tends to increase with the batch size due the lower number of rollout steps during training, which are a direct consequence of the higher memory footprint required for Curr training.
@@
Figure 9 provides additional samples of the trajectories presented in Figure 3 for the training with Pfwd. From Figure 9 we observe that models trained with the unregularized objectives remain highly dissipative despite using different random seeds. For this configuration of parameters we were able obtain only one stable model without the regularization among the random seeds chosen as shown in Figure 10. In this case, even though the trajectories are visually more realistic, if we consider the metrics used for evaluation Table 3 shows that the models trained with the regularization still provide better statistics.
@@
We point out that these highly dissipative models are also presented as we increase the batch size, for example Figures 11 and 12, show the same phenomenon for trajectories learned without regularization for batch sizes $256$ and $512$.
Refer to caption Refer to caption Refer to caption (a)
Refer to caption Refer to caption Refer to caption (b)
Refer to caption Refer to caption Refer to caption (c)
@@
Figure 8: Evolution of the cosine similiraty with time trajectories trained with both regularized and unregularized objectives for different batch sizes (a) $32$, (b) $64$ and (c) $128$. The solid line is the median among $160$ runs, and the shaded regions corresponds second and third quartile. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate= 5e-4).
@@
Figure 9: Further samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization as shown in Figure 3. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = $128$ and learning rate= $45e-$ 4).
@@
Figure 10: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization. In this case we show samples of the random seed with the best results for the unregularized training. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate= 5e-4).
@@
Figure 11: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization with a batch size of $256$ for the same random seed. ($\lambda_{1}=0$, $\lambda_{2}=100$, and learning rate= 5e-4).
@@
Figure 12: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization with a batch size of $512$ for the same random seed. ($\lambda_{1}=0$, $\lambda_{2}=100$, and learning rate= 5e-4).
@@
Table 3: Metrics of the best model using the unregularized pushforward loss versus the average of he models trained using the regularized pushforward loss, for batch size = 128 and learning rate = 5e-5.
MELR↓ (×$\times 10^{-2}$) MELRw↓ (×$\times 10^{-2}$) covRMSE↓ (×$\times 10^{-2}$)
Reg 2.34 0.588 7.89
@@
Appendix G Assets
G.1 Software and Libraries
In Table 4, we list relevant open-source software (and corresponding licenses) used in this work:
Table 4: Open source libraries (and corresponding licenses) used in this work.
@@ G.1 Software and Libraries
Pandas (pandas development team, 2020) BSD 3-Clause “New" or “Revised"
@@
Seaborn (Waskom, 2021) BSD 3-Clause “New" or “Revised"

### v2

@@ Appendix F Further Results
F.1 Lorenz 63
Figure 4 depicts the attractor for the Lorenz 63 system for an ensemble of trajectories (with randomly chosen initial conditions close to the attractor) at time $t=400$. We observe that DySLIM regularization provides some extra symmetry to the attractor, as evidenced by a better defined right-wing.
In addition, we provide several other metrics to showcase the advantage of our methodology. Figure 5 shows the behavior of the MMD metric for much longer horizons to the ones used for training. We can observe that DySLIM outperforms all the baselines. We also, studied the Wasserstein-1 distance of different features involved in the ground truth dynamics shown in Equation 32. In this case, Figure 6 shows that the distribution of each of the components is better captured in the models trained with DySLIM.
F.2 Kuramoto-Sivashinsky
In addition to the results shown in the main text, Figure 7 shows the improved stability from DySLIM by comparing the distribution of first order (${\bm{u}}_{x}$) and second order (${\bm{u}}_{xx}$) spatial derivatives for ground truth and predicted trajectories, which are relevant quantities that appear in the PDE that defines this system in Equation 33. We use finite difference methods to compute ${\bm{u}}_{x}$ and ${\bm{u}}_{xx}$ at each point in time for each trajectory in the test set. At each point in time, we thus have a distribution for these derivatives across test set trajectories and spatial grid. We use the Wasserstein-1 distance to compare ground truth and predicted distributions and find that models trained with regularized objectives better match the distribution of ground truth spatial derivatives. In summary, Figure 7 shows that by regularizing the loss, we obtain a closer distribution on the derivatives than when using the unregularized loss. In fact, for some cases of the curriculum training, the Wasserstein-1 matrix explodes as the trajectories are highly unstable.
Figure 7: Wasserstein-1 distance ($\downarrow$) on distribution of first order (Top) and second order (Bottom) spatial derivatives across time. Values greater than 3 are not shown on the plot. Each line corresponds to one of five random training seeds with bolded lines indicating median values (excluding trajectories that produce NaN values). The solid black line corresponds to statistics calculated from the distribution of initial conditions.
@@ F.3 Kolmogorov Flow
In this section, we provide ablation results and additional trajectory samples.
Table 5 compiles additional results for the model trained with and without regularization following the description in Appendix E.3.1. This table shows that using DySLIM either improves or roughly maintains values for all metrics. We point out that for curriculum training, the error tends to increase with batch size due to the lower number of rollout steps during training, which is a direct consequence of the higher memory footprint required for curriculum training.
@@
Table 6: Kolmogorov flow: metrics for pushforward training for $40$ time steps. ($\lambda_{1}=0$, $\lambda_{2}=100$). Boldface numbers indicate that the metric is improved by our regularization.
Batch size Learning rate
MELR ($\downarrow$)
(×$\times 10^{-2}$)
MELRw ($\downarrow$)
(×$\times 10^{-2}$)
covRMSE ($\downarrow$)
(×$\times 10^{-2}$)
Wass1 ($\downarrow$)
(×$\times 10^{-2}$)
TCM ($\downarrow$)
(×$\times 10^{-2}$)
Base DySLIM Base DySLIM Base DySLIM Base DySLIM Base DySLIM
Pushforward 128 5e-4 62.22 8.51 24.01 1.65 42.6 8.07 175 11.34 161 4.3
@@
Figure 9 provides additional samples of the trajectories presented in Figure 3 for the training with the pushforward objective. From Figure 9, we observe that models trained with the unregularized objectives remain highly dissipative despite using different random seeds. For this configuration of parameters, we were able obtain only one stable model trained without the regularization among the random seeds, which we present in Figure 10. In this case, even though the trajectories are visually more realistic, if we consider the metrics used for evaluation, Table 7 shows that the models trained with the DySLIM regularization still provide better statistics.
@@
We point out that these highly dissipative models are also present as we increase the batch size. For example, Figures 11 and 12 show the same phenomenon for trajectories learned without regularization for batch sizes of $256$ and $512$.
In addition, Table 6 shows the statistics of models trained with much longer time-horizons ($40$ time steps instead of $10$) for the Kolmogorov flow. The baseline in this case becomes completely uninformative, whereas the regularized version still provides models with reasonable statistics.
@@
Figure 8: Evolution of the cosine similarity ($\uparrow$) overt time for trajectories trained with both regularized and unregularized objectives for different batch sizes of (a) 32, (b) 64, and (c) 128. The solid line is the median among 160 runs, and the shaded regions corresponds second and third quartile. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate= $5\mathrm{e}^{-4}$).
@@
Figure 9: Additional samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained using the pushforward objective, with and without regularization. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = $128$ and learning rate= $5\mathrm{e}^{-4}$).
@@
Figure 10: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization. In this case, we show samples of the random seed with the best results for the unregularized training. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate= 5e-4).
@@
Figure 11: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization with a batch size of $256$ for the same random seed. ($\lambda_{1}=0$, $\lambda_{2}=100$, and learning rate = $5\mathrm{e}^{-4}$).
@@
Figure 12: Samples of reference trajectories of Navier Stokes system with Kolmogorov forcing and predicted trajectory generated with models trained the pushforward loss, with and without regularization with a batch size of $512$ for the same random seed. ($\lambda_{1}=0$, $\lambda_{2}=100$, and learning rate = $5\mathrm{e}^{-4}$).
@@
Table 7: Metrics of the best model using the unregularized pushforward loss versus the average of the models trained using the regularized pushforward loss, for batch size = 128 and learning rate = 5e-5.
MELR ($\downarrow$) (×$\times 10^{-2}$) MELRw ($\downarrow$) (×$\times 10^{-2}$) covRMSE ($\downarrow$) (×$\times 10^{-2}$)
@@
 +DySLIM 2.34 0.588 7.89
Appendix G Sinkhorn Divergence for Measure Matching
[…]
Lorenz 63 Ablation
From Figures 13, 14, and 15, we can observe that using SD to regularize the objectives provides some benefits. For the short term behavior in Figure 13, we see that the gains are very similar between using SD or MMD in place of D. For the measure matching metrics, we see that both measure-matching metrics stabilize the trajectories. This can be further be seen in Figures 16 and 17, which show that SD does stabilize some of the summary metrics. We point out that even though the performance is competitive, using MMD still seems to have an edge, albeit fairly small.
Kuramoto–Sivashinsky Ablation
In Figure 18, we find evidence that using SD as the regularizer for the KS system does improve stability, but performance with respect to using MMD starts to deteriorate. In Figure 18 (a), we see that there is indeed a better performance when using MMD as opposed to SD. In Figure 18 (b) we find that even though the SD regularization helps, many of the trajectories still diverge particularly for curriculum training.
The conclusion from this ablation is that while SD is compatible with our DySLIM framework and does provide increased stability, we find that using MMD in the regularizer leads to superior performance, especially for systems with larger dimension, e.g., KS.
Figure 13: Cosine similarity metric ($\uparrow$) for test trajectories at various rollout times of the Lorenz 63 system using MMD and SD for measure matching during training. Each line corresponds to the mean over trajectories of each of five random training seeds, with bold lines indicating median values.
Figure 14: Sinkhorn Divergence (SD; $\downarrow$) between trajectories at various rollout times of the Lorenz 63 system. (a) Values when using MMD for measure matching at training. (b) Values when using SD for measure matching at training. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Figure 15: Maximum Mean Discrepancy (MMD; $\downarrow$) between trajectories at various rollout times of the Lorenz 63 system. (a) Values when using MMD for measure matching at training. (b) Values when using SD for measure matching at training. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Figure 16: Wasserstein-1 metric ($\downarrow$) for the $xz$ values between trajectories at various rollout times of the Lorenz 63 system. (a) Values when using MMD for measure matching at training. (b) Values when using SD for measure matching at training. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Figure 17: Wasserstein-1 metric ($\downarrow$) for the $xy$ coordinates between trajectories at various rollout times of the Lorenz 63 system. (a) Values when using MMD for measure matching at training. (b) Values when using SD for measure matching at training. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Figure 18: Metrics for KS system. (a) Cosine Similarity ($\uparrow$) between the different trajectories. Each line corresponds to the mean over trajectories of each of five random training seeds, with bold lines indicating median values. (b) Sinkhorn Divergence (SD; $\downarrow$) between trajectories at various rollout times when using MMD for measure matching at training. (c) SD between trajectories at various rollout times when using SD for measure matching at training. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Appendix H Assets
H.1 Software and Libraries
In Table 8, we list relevant open-source software, and corresponding licenses, used in this work:
Table 8: Open source libraries used in this work, with corresponding licenses.
@@ H.1 Software and Libraries
Pandas (pandas development team, 2020) BSD 3-Clause “New” or “Revised”
@@
Seaborn (Waskom, 2021) BSD 3-Clause “New” or “Revised”